OpenStudy (anonymous):

can someone explain to me how to do this? don't just give me an answer, i honestly need some steps lol Which of the following represents the solution of x - 8 < -2 in set notation?

OpenStudy (mathstudent55):

To solve the inequality you need to get x by itself on the left side. Since 8 is being subtracted from x, you must add 8 to have x by itself. In an inequality, you must do the same operation to both sides, so add 8 to both sides of the inequality. What do you get? x - 8 < -2 + 8 +8 --------------

OpenStudy (anonymous):

ok, when adding to a negative number...do you add, or subtract? i can't remember the law of the whole...negative and positive numbers thing D:

OpenStudy (mathstudent55):

Here are the rules of addition of numbers. 1. To add two positive numbers, just add them as you've been adding since elementary school. Examples: 5 + 3 = 8; 1 + 2 = 3 2. To ad two negative numbers, just forget the signs, add the numbers as if they were both positive, then the answer is negative. Examples: -2 + (-5): Think of 2 + 5 = 7, but the answer is negative, so -2 + (-5) = -7 -23 + (-32): Think of 23 + 32 = 55. The answer is negative, so the answer is -55. 3. To add a positive and a negative number. Take the absolute value of both numbers. That means think of both numbers as being positive. Subtract the smaller number from the larger number. The answer has rthe sign of the larger number. Examples: -6 + 8. The absolute values are 6 and 8. Subtract the smaller number from the larger number, 8 - 6 = 2. The sign is the same as the larger original number. 8 is larger than 6 and is positive, so the answer is positive, so -6 + 8 = 2 9 + (-12). Take the absolute values: 12 and 9. Subtract: 12 - 9 = 3. Since 12 is the larger original number and is negative, the answer is negative. 9 + (-12) = -3.